Find the x-intercept for y=(6x+1)(2x-3)
Answer by jim_thompson5910(35256)

Plug in y = 0 and solve for x. You'll get 2 solutions.

y=(6x+1)(2x-3)
0=(6x+1)(2x-3)
(6x+1)(2x-3) = 0
6x+1 = 0 or 2x-3 = 0
6x = -1 or 2x = 3
x = -1/6 or x = 3/2

The two solutions are x = -1/6 or x = 3/2